# PolaRiS: Scalable Real-to-Sim Evaluations for Generalist Robot Policies

### v1

@@ 1 Introduction
Recently, generalist robotic policies [1, 2, 3, 4, 5, 6, 7, 8], trained on diverse real-robot datasets, have shown an impressive degree of generalization to novel scenarios. However, with the advent of generalist policies comes a critical demand for large-scale empirical validation, since absolute guarantees are challenging to provide. To enable generalization, evaluations must be conducted across a large variety of tasks, environments and rollouts. While performing evaluations directly on real-world robots is the gold standard, this quickly becomes tedious, hard to control systematically, and impractical for individual groups to perform due to cost and labor constraints. Therefore, a more scalable paradigm for generalist robot policy evaluation becomes an emerging demand.
@@
Simulation provides an appealing alternative—cheap, safe, fast, and unburdened by physical hardware, making large-scale policy evaluation far more feasible. However, despite considerable efforts [9, 10, 11, 12, 13, 14, 15, 16], these platforms have not been well suited for evaluating the aforementioned generalist policies. To support a wide range of scenarios required to test generality, existing robot-simulation benchmarks typically sacrifice realism—including dynamics, visuals, and scene geometry—and thus fail to capture the visual complexity of the physical world [9, 17, 14, 11, 12]. Our experiments show that there remains a poor correlation between simulation and real world policy performance of state-of-the-art generalist policies. Moreover, as generalist policies get deployed across diverse environments, roboticists are often interested not only in average model performance across environments, but also in understanding a policy’s performance within a specific target environment, which necessitates a tailored simulation of that scene. Constructing such an environment typically requires considerable manual effort. Effective simulated evaluation of generalist policies requires a new set of tools that allow us to rapidly construct high-fidelity targeted simulation environments, reflecting the diversity of real-world environments.
@@
To this end, we introduce PolaRiS, a scalable real-to-sim framework for evaluating generalist robot policies. PolaRiS leverages neural reconstruction such as Gaussian splatting [18, 19] and 3D object generation [20] to construct simulated evaluation environments that closely match their real-world counterparts with high visual and geometric fidelity (Figure˜1). This enables advanced generalist policies to perform various tasks in PolaRiS simulations out of the box. Additionally, our experiments show that lightweight finetuning of these policies with a small dataset of demonstrations in simulation can significantly boost real-to-sim correlation. Importantly, unlike existing simulation benchmarks, this simulation data is used only to align visual representations instead of teaching new tasks, and therefore does not have to be collected in the target environment or for the target task. As a result, once finetuned, policies can be evaluated zero-shot in unseen PolaRiS environments, facilitating rapid development of simulated evaluations that are customized for individual deployments yet shared broadly to assess generalist competence.
@@
In our experiments across diverse real-world environments using the DROID platform [21], PolaRiS evaluations match the real-world performance of state-of-the-art generalist policies out of the box with an average Pearson correlation of $r=0.9$. PolaRiS evaluations also strongly correlate ($r=0.98$) with policy scores in RoboArena [22], the most comprehensive real-world benchmark for generalist robot policies to date. As such, our experiments suggest that PolaRiS takes a significant step towards rapid, scalable, simulated evaluation of generalist policies. We open-source all tools required to create, use, and share PolaRiS evaluation environments, to foster the community-driven development of a shared suite of realistic simulation evaluation environments.
@@ 2 Related Work
Robot Policy Evaluation: Accurately evaluating the performance of robot policies is important for guiding research, yet faces challenges in practice: from reproducing initial environment conditions to robot wear and tear, and the tedium of supervising many hours of evaluations. Nevertheless, numerous real-robot benchmarks have been proposed over the years [23, 24, 25, 22]. With the rise of increasingly general policies that can perform many tasks across many environments [2, 1, 5, 6, 26], the challenges for performing comprehensive real-world evaluations are exacerbated. Simulated evaluations pose a promising alternative: initial conditions in simulation are perfectly repeatable, simulated robots do not wear out, and evaluations can be rapidly scaled through parallelization. As a result, many simulated robot benchmarks have been proposed [9, 11, 12, 13, 17]. These benchmarks are invaluable for algorithmic development. However, due to large visual and physical domain gaps between these simulated environments and the real world, they struggle to evaluate the performance of generalist policies, trained on real-world data. We instead design simulated environments that are capable of evaluating real-robot, generalist policies. They share the aforementioned benefits of simulated evaluations with respect to usability, reproducibility, and scalability, but are also reflective of the real-world performance of the evaluated generalist policies. More recently, advances in world models have enabled the use of action-conditioned video models as an alternative tool for robot policy evaluation [27, 28, 29, 30]. Like “classic” simulation, video model evaluations can be run at scale. Yet, our experiments demonstrate that today’s (open) models struggle to reliably model physical interactions, and thus struggle to estimate policy performance. In contrast, our method leverages an accurate physics simulator together with high-fidelity neural rendering, making policy evaluation more faithful to real-world dynamics, reliable, efficient, and cost-effective.
@@
Real-to-sim Policy Evaluation: A few recent works have recently explored whether increased visual and physical realism can improve the ability of simulated environments to be predictive of real-world policy performance: [31, 32] demonstrated that navigation policies can be evaluated in realistic reconstructions of indoor environments, and [33] demonstrated the same for autonomous driving policies. In robot manipulation, where policies closely interact with their environment, the requirements for physical and visual realism of simulated evaluation environments increase: SIMPLER [9] first showed that simulation can be used for evaluation of real-robot manipulation policies, but it fails to deliver strong real-world performance correlation for recent generalist policies [8]. Additionally, the “green-screening” approach used by [9] and follow-up works [34], which copy-pastes background pixels from real-robot images, does not support evaluation for robot setups with moving cameras. Unfortunately, this excludes most generalist policies [5, 26, 2, 1], which feature moving wrist cameras attached to the robot’s end-effector. Our work uses 3D Gaussian splatting techniques to create high-fidelity simulation environments. The resulting environments show strong real-to-sim performance correlation and support the evaluations of policies with wrist cameras.
@@
Real-to-sim Environment Creation: A number of works have developed methods for digitizing real-world environments. Some approaches focus solely on room layout, and fill the scene by sampling from a large 3D object database [35], often lacking in visual realism. For higher visual quality, early approaches relied on specialized and expensive sensory equipment [36, 37], and more recently, learned models have been used to bypass such requirements and instead infer 3D models of environments from raw image or video inputs [38, 32]. In this context, Gaussian splatting has become a popular tool for inferring scene geometry and appearance from raw sensory observations [19, 39, 18, 40]. Multiple works have explored the use of Gaussian splatting techniques in robotics [41, 42, 43, 44, 45, 46]. While these works have shown that splatting can produce environments that can be used for augmenting the training of robot policies, in this work we investigate the utility of real-to-sim techniques in the context of robot policy evaluation. Most closely related to our work, [47, 48] used simulated environments created via Gaussian splatting for policy evaluation, but focused on evaluating narrow, single-task policies in environments seen in the training data. In contrast, comprehensive evaluation of modern generalist policies requires evaluations across a large number of tasks in unseen environments [22]. Our work is the first to develop a pipeline that enables evaluation of real-robot policies in unseen simulated environments, with strong performance correlation to real-world evaluations.
@@ 3 Preliminaries
For each policy $\pi_{i}\in\Pi$, we can define a metric of expected real-world performance (e.g., success rate) as $R_{i}=\mathbb{E}_{e\sim\mathcal{E}_{\text{real}}}\big[\,r(e,\pi_{i})\,\big]$, where $r(e,\pi_{i})$ denotes the episodic performance when executing policy $\pi_{i}$ in environment $e$. Analogously, we define the policy’s performance in simulation as $R_{\mathcal{S},i}=\mathbb{E}_{e\sim\mathcal{E}_{\mathcal{S}}}\big[\,r(e,\pi_{i})\,\big].$ The goal of simulation-based evaluation is to design simulation environments $\mathcal{E}_{\mathcal{S}}$ such that the ranking and relative performance of policies in simulation closely approximate their real-world behavior. We follow prior work [9] to quantify the faithfulness of policy evaluation using two metrics.
@@
2D Gaussian Splatting for Scene Reconstruction: To faithfully recover the scene geometry, we resort to recent neural reconstruction tools [19, 18, 49, 50, 39]. To achieve the best trade-off between high visual fidelity and accurate mesh geometry, we leverage 2D Gaussian Splatting (2DGS) [18], which represents a scene as a collection of oriented planar Gaussian disks embedded in 3D space, enabling improved geometric fidelity compared to other representations. A 2D Gaussian splat is represented as a tuple $g_{j}=(\bm{\mu}_{j},\;\mathbf{R}_{j},\;\mathbf{S}_{j},\;\mathbf{c}_{j},\;\alpha_{j}),$ where $\bm{\mu}_{j}\in\mathbb{R}^{3}$ is the disk center, $\mathbf{R}_{j}\in\mathrm{SO}(3)$ encodes its orientation, $\mathbf{S}_{j}\in\mathbb{R}^{3\times 3}$ defines its covariance in the disk plane, and $(\mathbf{c}_{j},\alpha_{j})$ denote its color and opacity. A full scene is represented as a collection of these Gaussian splats $\mathcal{D}=\{g_{j}\}_{j=1}^{M}$, optimized so that a differentiable renderer reproduces the input views on aggregate. Once optimized, the resulting representation supports both real-time photorealistic rendering and accurate mesh extraction, making it well suited for physics-based simulation of robotic policies. Additional implementation details are provided in Appendix A.1.
Generalist Robot Policies: While generalist policies is generally an ill-defined term, in this work we specifically use it to refer to a class of language-conditioned multi-task robot policies known as Vision-Language-Action models (VLAs) [5, 2, 1, 26]. VLAs are built on top of transformer based Vision-Language Models (VLMs), fine-tuned to produce robot actions conditioned on visual observations and textual instructions. We denote a generalist policy as $\pi_{\theta}:(\mathcal{O}_{\text{vis}},\mathcal{O}_{\text{lang}})\rightarrow\mathcal{A}$, where $\mathcal{O}_{\text{vis}}$ is the space of observations, $\mathcal{O}_{\text{lang}}$ is the space of language instructions, and $\mathcal{A}$ is the action space.
@@ 4.1 Environment Generation with PolaRiS
Environment Creation via Reconstruction: We start by capturing a short (2–5 minutes) monocular video recording of the scene. We then run a reconstruction pipeline based on 2DGS [18] to reconstruct scene visuals and geometry. To facilitate compositional scene construction, we record separate video sequences for (i) the static background scene and (ii) the robot. This enables training of distinct Gaussian splat models for each component, which can later be composed and articulated in Euclidean space. During capture, the user slowly moves the camera to avoid motion blur while ensuring full coverage of all visible surfaces. To assist in automatically resolving global orientation and metric scale, a ChArUco board is placed in the environment (see Figure˜2). This data acquisition process for a scene typically requires only 5–10 minutes. Following data capture, we estimate the camera parameters for all frames using COLMAP [51] and train a 2D Gaussian Splatting model $\mathcal{D}=\{g_{j}\}_{j=1}^{M};g_{j}=(\bm{\mu}_{j},\;\mathbf{R}_{j},\;\mathbf{S}_{j},\;\mathbf{c}_{j},\;\alpha_{j})$ for each recording by optimizing a standard photometric reconstruction loss as described in Appendix A.1.
To enable closed-loop robotic evaluation in a physics-based simulator, we require the geometric and photometric components of the reconstructed scene. We obtain a Truncated Signed Distance Field (TSDF) using a combination of rasterizing the splats and volumetric depth fusion [52]. An explicit mesh representation $\mathcal{M}$ is then extracted from the TSDF using marching cubes [53]. The resulting meshes are imported into IsaacSim [54] to enable realistic contact dynamics and object interactions, while the underlying Gaussian splats $\mathcal{D}$ provides the corresponding photometric renders for high-fidelity visual affects. The environment creation framework can be fully automated and typically completes within 30 minutes on an NVIDIA RTX 4090. Please refer to Appendix B for implementation details.
@@
Object Creation with Generative Models: We find that directly extracting high-quality, watertight assets from real-world scenes is a time-consuming and tedious process that often requires depth-enabled hardware and precise registration tools. Although PolaRiS is not tied to any specific object reconstruction methods, we employ the TRELLIS 3D image-to-3D generative model [20] for simplicity and flexibility. TRELLIS introduces a structured representation that enables mapping from a single or a few multiview 2D images to diverse 3D representations (such as Gaussian splats, radiance fields, and meshes) with high fidelity, while recovering unobservable object regions. With multiview images captured with a standard camera as inputs, we segment the object using SAM2 [55] and feed the segmented views into TRELLIS to obtain a structured latent representation. This latent is decoded into a Gaussian splat for appearance and an untextured mesh for geometry, after which appearance is baked onto the mesh to produce a fully textured object for simulation.
@@
Composing simulated evaluation scenes. The final step is to assemble these components into an interactive evaluation scene using our Scene Composition GUI (Figure˜3). Since all components are represented as Gaussian primitives, the environment, robot and objects can be composed and transformed in Euclidean space. Moreover, we set standard physics parameters for contact dynamics using default values from IsaacSim, and use manufacturer-provided defaults for the robot arm. Object mass parameters can be estimated by the users. Once the scene is assembled, it can be exported as a USD file for evaluation. Alongside scene composition, the user also defines a success criterion for scoring evaluation rollouts – e.g., whether a particular object is placed within a certain receptacle. The total human effort required for creating a new evaluation environment is typically less than 20 minutes, and the total wall time for creating a new evaluation environment, including training time of the Gaussian splats, is typically less than one hour. This is more scalable than prior simulated evaluation works [9], which require hand-defined construction of the environment geometry and extensive tuning of lighting and textures.
An important facet of the constructed simulation is the ability to provide policies with rich sensor inputs. As opposed to prior work [9, 34] which copied image backgrounds in 2D via green-screen, PolaRiS reconstructs the entire scene in 3D, providing full 3D rendering and novel view synthesis. This allows us to easily render third person views from any angle and provide wrist camera inputs to the policy. This opens up a new class of simulation evaluations for generalist policies that prior real-to-sim evaluation frameworks that rely on background inpainting [9, 34] were unable to support.

### v2

@@ 1 Introduction
Recently, generalist robotic policies [blackpi0, pi05, o2024open, team2025gemini, geminirobotics, lbmtri2025, octo_2023, kim24openvla], trained on diverse real-robot datasets, have shown an impressive degree of generalization to novel scenarios. However, with the advent of generalist policies comes a critical demand for large-scale empirical validation, since absolute guarantees are challenging to provide. To enable generalization, evaluations must be conducted across a large variety of tasks, environments and rollouts. While performing evaluations directly on real-world robots is the gold standard, this quickly becomes tedious, hard to control systematically, and impractical for individual groups to perform due to cost and labor constraints. Therefore, a more scalable paradigm for generalist robot policy evaluation becomes an emerging demand.
@@
Simulation provides an appealing alternative—cheap, safe, fast, and unburdened by physical hardware, making large-scale policy evaluation far more feasible. However, despite considerable efforts [li24simpler, katara24gensim, liu23libero, nasiriany24robocasa, james20rlbench, mees22calvin, yu2020meta, li2024behavior], these platforms have not been well suited for evaluating the aforementioned generalist policies. To support a wide range of scenarios required to test generality, existing robot-simulation benchmarks typically sacrifice realism—including dynamics, visuals, and scene geometry—and thus fail to capture the visual complexity of the physical world [li24simpler, yu19metaworld, mees22calvin, liu23libero, nasiriany24robocasa]. Our experiments show that there remains a poor correlation between simulation and real world policy performance of state-of-the-art generalist policies. Moreover, as generalist policies get deployed across diverse environments, roboticists are often interested not only in average model performance across environments, but also in understanding a policy’s performance within a specific target environment, which necessitates a tailored simulation of that scene. Constructing such an environment typically requires considerable manual effort. Effective simulated evaluation of generalist policies requires a new set of tools that allow us to rapidly construct high-fidelity targeted simulation environments, reflecting the diversity of real-world environments.
@@
To this end, we introduce PolaRiS, a scalable real-to-sim framework for evaluating generalist robot policies. PolaRiS leverages neural reconstruction such as Gaussian splatting [2dgs, 3dgs] and 3D object generation [xiang2025structured] to construct simulated evaluation environments that closely match their real-world counterparts with high visual and geometric fidelity (Figure˜1). This enables advanced generalist policies to perform various tasks in PolaRiS simulations out of the box. Additionally, our experiments show that lightweight finetuning of these policies with a small dataset of demonstrations in simulation can significantly boost real-to-sim correlation. Importantly, unlike existing simulation benchmarks, this simulation data is used only to align visual representations instead of teaching new tasks, and therefore does not have to be collected in the target environment or for the target task. As a result, once finetuned, policies can be evaluated zero-shot in unseen PolaRiS environments, facilitating rapid development of simulated evaluations that are customized for individual deployments yet shared broadly to assess generalist competence.
@@
In our experiments across diverse real-world environments using the DROID platform [khazatsky24droid], PolaRiS evaluations match the real-world performance of state-of-the-art generalist policies out of the box with an average Pearson correlation of $r=0.9$. PolaRiS evaluations also strongly correlate ($r=0.98$) with policy scores in RoboArena [atreya2025roboarena], the most comprehensive real-world benchmark for generalist robot policies to date. As such, our experiments suggest that PolaRiS takes a significant step towards rapid, scalable, simulated evaluation of generalist policies. We open-source all tools required to create, use, and share PolaRiS evaluation environments, to foster the community-driven development of a shared suite of realistic simulation evaluation environments.
@@ 2 Related Work
Robot Policy Evaluation: Accurately evaluating the performance of robot policies is important for guiding research, yet faces challenges in practice: from reproducing initial environment conditions to robot wear and tear, and the tedium of supervising many hours of evaluations. Nevertheless, numerous real-robot benchmarks have been proposed over the years [calli15ycb, ahn19robel, zhou25autoeval, atreya2025roboarena]. With the rise of increasingly general policies that can perform many tasks across many environments [pi05, blackpi0, geminirobotics, lbmtri2025, nvidia2025gr00tn1openfoundation], the challenges for performing comprehensive real-world evaluations are exacerbated. Simulated evaluations pose a promising alternative: initial conditions in simulation are perfectly repeatable, simulated robots do not wear out, and evaluations can be rapidly scaled through parallelization. As a result, many simulated robot benchmarks have been proposed [li24simpler, liu23libero, nasiriany24robocasa, james20rlbench, yu19metaworld]. These benchmarks are invaluable for algorithmic development. However, due to large visual and physical domain gaps between these simulated environments and the real world, they struggle to evaluate the performance of generalist policies, trained on real-world data. We instead design simulated environments that are capable of evaluating real-robot, generalist policies. They share the aforementioned benefits of simulated evaluations with respect to usability, reproducibility, and scalability, but are also reflective of the real-world performance of the evaluated generalist policies. More recently, advances in world models have enabled the use of action-conditioned video models as an alternative tool for robot policy evaluation [quevedo2025evaluating, tseng2025scalable, guo2025ctrl, veorobotics2025]. Like “classic” simulation, video model evaluations can be run at scale. Yet, our experiments demonstrate that today’s (open) models struggle to reliably model physical interactions, and thus struggle to estimate policy performance. In contrast, our method leverages an accurate physics simulator together with high-fidelity neural rendering, making policy evaluation more faithful to real-world dynamics, reliable, efficient, and cost-effective.
@@
Real-to-sim Policy Evaluation: A few recent works have recently explored whether increased visual and physical realism can improve the ability of simulated environments to be predictive of real-world policy performance: [chhablani2025embodiedsplatpersonalizedrealtosimtorealnavigation, escontrela2025gaussgymopensourcerealtosimframework] demonstrated that navigation policies can be evaluated in realistic reconstructions of indoor environments, and [yang2023unisim] demonstrated the same for autonomous driving policies. In robot manipulation, where policies closely interact with their environment, the requirements for physical and visual realism of simulated evaluation environments increase: SIMPLER [li24simpler] first showed that simulation can be used for evaluation of real-robot manipulation policies, but it fails to deliver strong real-world performance correlation for recent generalist policies [kim24openvla]. Additionally, the “green-screening” approach used by [li24simpler] and follow-up works [jangir2025robotarenainftyscalablerobot], which copy-pastes background pixels from real-robot images, does not support evaluation for robot setups with moving cameras. Unfortunately, this excludes most generalist policies [geminirobotics, nvidia2025gr00tn1openfoundation, pi05, blackpi0], which feature moving wrist cameras attached to the robot’s end-effector. Our work uses 3D Gaussian splatting techniques to create high-fidelity simulation environments. The resulting environments show strong real-to-sim performance correlation and support the evaluations of policies with wrist cameras.
@@
Real-to-sim Environment Creation: A number of works have developed methods for digitizing real-world environments. Some approaches focus solely on room layout, and fill the scene by sampling from a large 3D object database [deitke2023phone2proc], often lacking in visual realism. For higher visual quality, early approaches relied on specialized and expensive sensory equipment [savva2019habitat, xia2018gibson], and more recently, learned models have been used to bypass such requirements and instead infer 3D models of environments from raw image or video inputs [torne2024rialto, escontrela2025gaussgymopensourcerealtosimframework]. In this context, Gaussian splatting has become a popular tool for inferring scene geometry and appearance from raw sensory observations [3dgs, chen2025survey3dgaussiansplatting, 2dgs, irshad2024neuralfieldsroboticssurvey]. Multiple works have explored the use of Gaussian splatting techniques in robotics [xia2025drawerdigitalreconstructionarticulation, yu2025pogs, chakra25realissim, yu2024language, qureshi2025splatsim, dan2025xsim, yu2025real2render2realscalingrobotdata]. While these works have shown that splatting can produce environments that can be used for augmenting the training of robot policies, in this work we investigate the utility of real-to-sim techniques in the context of robot policy evaluation. Most closely related to our work, [jiang2025gsworld, zhang2025real] used simulated environments created via Gaussian splatting for policy evaluation, but focused on evaluating narrow, single-task policies in environments seen in the training data. In contrast, comprehensive evaluation of modern generalist policies requires evaluations across a large number of tasks in unseen environments [atreya2025roboarena]. Our work is the first to develop a pipeline that enables evaluation of real-robot policies in unseen simulated environments, with strong performance correlation to real-world evaluations.
@@ 3 Preliminaries
For each policy $\pi_{i}\in\Pi$, we can define a metric of expected real-world performance (e.g., success rate) as $R_{i}=\mathbb{E}_{e\sim\mathcal{E}_{\text{real}}}\big[\,r(e,\pi_{i})\,\big]$, where $r(e,\pi_{i})$ denotes the episodic performance when executing policy $\pi_{i}$ in environment $e$. Analogously, we define the policy’s performance in simulation as $R_{\mathcal{S},i}=\mathbb{E}_{e\sim\mathcal{E}_{\mathcal{S}}}\big[\,r(e,\pi_{i})\,\big].$ The goal of simulation-based evaluation is to design simulation environments $\mathcal{E}_{\mathcal{S}}$ such that the ranking and relative performance of policies in simulation closely approximate their real-world behavior. We follow prior work [li24simpler] to quantify the faithfulness of policy evaluation using two metrics.
@@
2D Gaussian Splatting for Scene Reconstruction: To faithfully recover the scene geometry, we resort to recent neural reconstruction tools [3dgs, 2dgs, mildenhall2020nerf, bakedsdf, chen2025survey3dgaussiansplatting]. To achieve the best trade-off between high visual fidelity and accurate mesh geometry, we leverage 2D Gaussian Splatting (2DGS) [2dgs], which represents a scene as a collection of oriented planar Gaussian disks embedded in 3D space, enabling improved geometric fidelity compared to other representations. A 2D Gaussian splat is represented as a tuple $g_{j}=(\bm{\mu}_{j},\;\mathbf{R}_{j},\;\mathbf{S}_{j},\;\mathbf{c}_{j},\;\alpha_{j}),$ where $\bm{\mu}_{j}\in\mathbb{R}^{3}$ is the disk center, $\mathbf{R}_{j}\in\mathrm{SO}(3)$ encodes its orientation, $\mathbf{S}_{j}\in\mathbb{R}^{3\times 3}$ defines its covariance in the disk plane, and $(\mathbf{c}_{j},\alpha_{j})$ denote its color and opacity. A full scene is represented as a collection of these Gaussian splats $\mathcal{D}=\{g_{j}\}_{j=1}^{M}$, optimized so that a differentiable renderer reproduces the input views on aggregate. Once optimized, the resulting representation supports both real-time photorealistic rendering and accurate mesh extraction, making it well suited for physics-based simulation of robotic policies. Additional implementation details are provided in Appendix A.1.
Generalist Robot Policies: While generalist policies is generally an ill-defined term, in this work we specifically use it to refer to a class of language-conditioned multi-task robot policies known as Vision-Language-Action models (VLAs) [geminirobotics, pi05, blackpi0, nvidia2025gr00tn1openfoundation]. VLAs are built on top of transformer based Vision-Language Models (VLMs), fine-tuned to produce robot actions conditioned on visual observations and textual instructions. We denote a generalist policy as $\pi_{\theta}:(\mathcal{O}_{\text{vis}},\mathcal{O}_{\text{lang}})\rightarrow\mathcal{A}$, where $\mathcal{O}_{\text{vis}}$ is the space of observations, $\mathcal{O}_{\text{lang}}$ is the space of language instructions, and $\mathcal{A}$ is the action space.
@@ 4.1 Environment Generation with PolaRiS
Environment Creation via Reconstruction: We start by capturing a short (2–5 minutes) monocular video recording of the scene. We then run a reconstruction pipeline based on 2DGS [2dgs] to reconstruct scene visuals and geometry. To facilitate compositional scene construction, we record separate video sequences for (i) the static background scene and (ii) the robot. This enables training of distinct Gaussian splat models for each component, which can later be composed and articulated in Euclidean space. During capture, the user slowly moves the camera to avoid motion blur while ensuring full coverage of all visible surfaces. To assist in automatically resolving global orientation and metric scale, a ChArUco board is placed in the environment (see Figure˜2). This data acquisition process for a scene typically requires only 5–10 minutes. Following data capture, we estimate the camera parameters for all frames using COLMAP [schonberger2016colmap] and train a 2D Gaussian Splatting model $\mathcal{D}=\{g_{j}\}_{j=1}^{M};g_{j}=(\bm{\mu}_{j},\;\mathbf{R}_{j},\;\mathbf{S}_{j},\;\mathbf{c}_{j},\;\alpha_{j})$ for each recording by optimizing a standard photometric reconstruction loss as described in Appendix A.1.
To enable closed-loop robotic evaluation in a physics-based simulator, we require the geometric and photometric components of the reconstructed scene. We obtain a Truncated Signed Distance Field (TSDF) using a combination of rasterizing the splats and volumetric depth fusion [curless1996tsdf]. An explicit mesh representation $\mathcal{M}$ is then extracted from the TSDF using marching cubes [LorensenCline87]. The resulting meshes are imported into IsaacSim [NVIDIA_Isaac_Sim] to enable realistic contact dynamics and object interactions, while the underlying Gaussian splats $\mathcal{D}$ provides the corresponding photometric renders for high-fidelity visual affects. The environment creation framework can be fully automated and typically completes within 30 minutes on an NVIDIA RTX 4090. Please refer to Appendix B for implementation details.
@@
Object Creation with Generative Models: We find that directly extracting high-quality, watertight assets from real-world scenes is a time-consuming and tedious process that often requires depth-enabled hardware and precise registration tools. Although PolaRiS is not tied to any specific object reconstruction methods, we employ the TRELLIS 3D image-to-3D generative model [xiang2025structured] for simplicity and flexibility. TRELLIS introduces a structured representation that enables mapping from a single or a few multiview 2D images to diverse 3D representations (such as Gaussian splats, radiance fields, and meshes) with high fidelity, while recovering unobservable object regions. With multiview images captured with a standard camera as inputs, we segment the object using SAM2 [ravi25sam2] and feed the segmented views into TRELLIS to obtain a structured latent representation. This latent is decoded into a Gaussian splat for appearance and an untextured mesh for geometry, after which appearance is baked onto the mesh to produce a fully textured object for simulation.
@@
Composing simulated evaluation scenes. The final step is to assemble these components into an interactive evaluation scene using our Scene Composition GUI (Figure˜3). Since all components are represented as Gaussian primitives, the environment, robot and objects can be composed and transformed in Euclidean space. Moreover, we set standard physics parameters for contact dynamics using default values from IsaacSim, and use manufacturer-provided defaults for the robot arm. Object mass parameters can be estimated by the users. Once the scene is assembled, it can be exported as a USD file for evaluation. Alongside scene composition, the user also defines a success criterion for scoring evaluation rollouts – e.g., whether a particular object is placed within a certain receptacle. The total human effort required for creating a new evaluation environment is typically less than 20 minutes, and the total wall time for creating a new evaluation environment, including training time of the Gaussian splats, is typically less than one hour. This is more scalable than prior simulated evaluation works [li24simpler], which require hand-defined construction of the environment geometry and extensive tuning of lighting and textures.
An important facet of the constructed simulation is the ability to provide policies with rich sensor inputs. As opposed to prior work [li24simpler, jangir2025robotarenainftyscalablerobot] which copied image backgrounds in 2D via green-screen, PolaRiS reconstructs the entire scene in 3D, providing full 3D rendering and novel view synthesis. This allows us to easily render third person views from any angle and provide wrist camera inputs to the policy. This opens up a new class of simulation evaluations for generalist policies that prior real-to-sim evaluation frameworks that rely on background inpainting [li24simpler, jangir2025robotarenainftyscalablerobot] were unable to support.